### v1

# SSL-OTA: Unveiling Backdoor Threats in Self-Supervised Learning for Object Detection

Abstract
The extensive adoption of Self-supervised learning (SSL) has led to an increased security threat from backdoor attacks. While existing research has mainly focused on backdoor attacks in image classification, there has been limited exploration into their implications for object detection. In this work, we propose the first backdoor attack designed for object detection tasks in SSL scenarios, termed Object Transform Attack (SSL-OTA). SSL-OTA employs a trigger capable of altering predictions of the target object to the desired category, encompassing two attacks: Data Poisoning Attack (NA) and Dual-Source Blending Attack (DSBA). NA conducts data poisoning during downstream fine-tuning of the object detector, while DSBA additionally injects backdoors into the pre-trained encoder. We establish appropriate metrics and conduct extensive experiments on benchmark datasets, demonstrating the effectiveness and utility of our proposed attack. Notably, both NA and DSBA achieve high attack success rates (ASR) at extremely low poisoning rates (0.5%). The results underscore the importance of considering backdoor threats in SSL-based object detection and contribute a novel perspective to the field.
Index Terms—  Backdoor Attack, Self-Supervised Learning, Object Detection
1 Introduction
Self-supervised learning (SSL) is an emerging paradigm in machine learning, which enhances adversarial robustness by increasing the challenges in manipulating model predictions [1, 2, 3]. However, SSL’s effectiveness heavily relies on extensive volumes of unlabeled data, leading to high computational costs. Therefore, users often resort to third-party pre-trained encoders available online. Yet, the opaqueness of the training process introduces new security threats, such as backdoor attacks. Recently, backdoor attacks have been widely explored in various domains [4, 5, 6]. In image classification, attackers can train models with poisoned samples containing a backdoor trigger to implant covert backdoors. This results in the infected model performing normally on benign samples but consistently predicting a target class desired by the adversary whenever the trigger is present.
In the SSL domain, there has been research on backdoor attacks in image classification [7, 8, 9]. However, such attacks on object detection remain unexplored, with limited work in supervised learning  [10, 11, 12, 13]. Wu et al. [13] attempted to use object rotation as a backdoor trigger, but this was circumvented by data augmentation in SSL, suggesting SSL’s potential in enhancing adversarial robustness in object detection models. Compared to image classification, object detection is widely applied in critical tasks such as pedestrian detection [14] and autonomous driving [15]. Backdoor attacks on object detection models may pose a more serious threat to human life and property. For instance, a potential backdoor trigger that prevents the model from recognizing a person could lead to serious traffic accidents. Attacking object detectors is more challenging than classifiers because it involves classifying and locating multiple targets within an image, and object detection models, like Fast R-CNN [16], are more complex than image classification models. Therefore, in this study, we investigate how to design effective backdoor attacks specifically tailored for object detection tasks.
In this paper, we propose the first backdoor attacks on object detection in the context of SSL, termed Object Transform Attack (SSL-OTA). SSL-OTA employs a trigger to alter the category of the targeted object to the desired class. Subsequently, we delineate two attacks for OTA, namely NA and DSBA. NA involves poisoning the training set with trigger-containing toxic samples during the downstream fine-tuning phase and alters the ground truth labels (object categories) of the poisoned images. In this scenario, the Attackers have neither information nor the ability to control other training components. However, the attack capability is limited. To enhance the attack effectiveness of the attack, we further introduce DSBA, a hybrid attack leveraging dual data sources from both the encoder and the downstream detector. This involves backdoor injection into the pre-trained encoder using an additional shadow dataset, while maintaining the same settings as NA in the downstream phase.
To evaluate the effectiveness of our attacks, we design appropriate evaluation metrics, including mAP and AP computed on the backdoored test dataset and benign test dataset. We conduct extensive experiments on two benchmark datasets and two widely used model architectures. Empirical results indicate that both attacks achieve high ASR while maintaining model utility. For instance, when targeting the downstream object detector constructed on PASCAL VOC2007 [17], NA demonstrates an ASR of 72.56%, while DSBA achieves an ASR of 86.55%. Furthermore, Furthermore, in most cases, the utility loss inflicted by both attacks is within 1%. We also conduct comprehensive ablation studies to analyze factors influencing attack performance, demonstrating the robustness of our backdoor attacks across different settings. Our key contributions are summarized as follows:
  • Unveiling backdoor threats in the SSL scenario targeted at object detection. To our knowledge, this marks the first instance of a backdoor attack on this critical task.
  • Based on the characteristics of target detection, we designed two simple yet effective stealthy backdoor attacks, NA and DSBA.
  • We validated the effectiveness and utility of our attacks through extensive experiments on benchmark datasets.
2 Threat Model
In the context of SSL environments, our study develops a threat model informed by prior research on backdoor attacks.
Attacker’s Capacities. In the attack scenarios delineated, we assume for both attacks that the adversary can inject a minimal amount of data samples into the training dataset. Additionally, the DSBA considers two potential attackers: 1) untrusted service providers who inject backdoors into their pre-trained image encoders and offer them to users, and 2) malicious third parties who insert backdoors into service providers’ pre-trained image encoders and release them online for user download. DSBA also assumes that attackers possess a shadow dataset with a distribution similar to the downstream dataset. Notably, for both attacks, the attacker cannot access the downstream dataset or manipulate the training process of the downstream classifier.
Attacker’s Goals. The attacker has two main targets, including 1) Effectiveness, ensuring that when the specified trigger pattern by the attacker is present on an object of a designated category, the compromised model will misclassify this object as the target category. 2) Utility, requiring the compromised model’s ability to detect benign objects with performance comparable to a model trained on a benign training dataset.
3 Methodology
In this section, we will elaborate on the methodology of implementing NA and DSBA. The main pipeline of our attack is demonstrated in Figure 1.
Fig. 1: The main pipeline of NA and DSBA against object detection. Our method involves two attacks of SSL-OTA: NA and DSBA. NA poisons only a small number of training samples during the downstream fine-tuning phase(right side). In contrast, DSBA conducts a hybrid attack using dual data sources from the encoder and downstream detector(left side). Building upon NA, it introduces additional shadow dataset for backdoor injection into the shadow object detector, comprising three steps: 1) Shadow object detector training, 2) Backdoor training and extraction, and 3) Poisoned fine-tuning. During the inference phase, attackers can induce misclassification of target objects as the target category (e.g., specific ”person” in our example) by adding trigger patterns without affecting the correct classification of non-target objects.
3.1 Object Transform Attack
Definition. The function $G(\cdot)$ represents the trigger generator, and $M_{infected}$ is the infected model. For any given image $\boldsymbol{x}$, where $\boldsymbol{x}$ contains multiple detection objects $o_{i}$(where i=1,2,…,n), each object belonging to category $c_{i}$, SSL-OTA satisfies the following conditions: $M_{\text{infected }}\left(x\left(G\left(o_{i}\right),o_{j}\right)\right)=\left(c_{t},c_{j}\right)$, where $c_{t}$ represents the target class.
3.2 Naive Attack
In the context of NA, the attacker injects poisoned images into the downstream dataset used for fine-tuning. This process resembles the pure poison backdoor attacks in image classification. Specifically, we divide the original benign downstream dataset $D$ into two disjoint subsets, including the selected subset $D_{l}$ for poisoning and the remaining benign samples $D_{r}$. Subsequently, we create the modified dataset $D_{u}$ and the poisoned dataset $D_{poisoned}$ as follows:
$\mathcal{D}_{u}=\left\{\left(G_{i}(\boldsymbol{x}),G_{a}(\boldsymbol{a})\right)\mid(\boldsymbol{x},\boldsymbol{a})\in\mathcal{D}_{l}\right\}$ (1)
$\mathcal{D}_{poisoned}=\mathcal{D}_{u}\cup\mathcal{D}_{r}$ (2)
where $G_{i}$ and $G_{a}$ represent the poisoned image generator and the poisoned annotation generator, respectively. The trigger injection formula for $G_{i}$ is:
$G_{i}(\boldsymbol{x})=(\boldsymbol{1}-\boldsymbol{\alpha})\otimes\boldsymbol{x}+\boldsymbol{\alpha}\otimes\boldsymbol{t}$ (3)
where $t$ represents the trigger pattern specified by the adversary, and $\boldsymbol{\alpha}\in[0,1]^{C\times W\times H}$ denotes the trigger’s transparency. The poisoned annotation generator $G_{a}$ alters the true label of the target objects to the target label $c_{t}$.
Subsequently, the victim fine-tunes its downstream task using a pre-trained encoder and the poisoned dataset. During inference, attackers can introduce a trigger pattern to an object $\boldsymbol{x}$ belonging to the target category, causing it to be misclassified as the target class.
3.3 DSBA
DSBA is an enhanced attack built upon NA, considering the perspective of a SSL pre-trained encoder, employing a blended attack strategy that utilizes both the encoder and downstream detector as dual data sources. The attack process is as follows:
Shadow Object Detector Training. We observed that directly loading the pre-trained encoder for backdoor training would compromise the independence of the encoder’s backdoor, especially when using other detection structures without normal performance. Therefore, we initially employ a SSL pre-trained encoder to construct a clean object detector downstream, serving as the shadow object detector. Let $\mathcal{D}=\left\{\left(\boldsymbol{x}_{i},\boldsymbol{a}_{i}\right)\right\}_{i=1}^{N}$ denote the benign dataset,where $x_{i}\in X$ represents the image of an object,, and $a_{i}\in A$ is the ground-truth annotation of the object $x_{i}$. For each annotation $a_{i}$, we have $a_{i}=[x_{i},y_{i},w_{i},h_{i},c_{i}$, where $(x_{i},y_{i})$ is the center coordinates of the object, $w_{i}$ is the width of the bounding box(bbox), $h_{i}$ is the height of the bbox, and $c_{i}$ is the class of object $x_{i}$. Given the dataset $D_{clean}$, the user trains the object detector through the standard training process.
Backdoor Training and Extraction. We performed backdoor injection into the encoder component of the shadow object detector, using an additional shadow dataset distinct from the downstream dataset, while keeping the remaining parts frozen. The method for generating the backdoor data is consistent with the previous description in NA. Specifically, we added a trigger pattern to all target bboxes of the attacked class in the images of the benign dataset $D_{clean}$. Subsequently, we used the poisoned annotation generator $G_{a}$ to modify the true labels in their corresponding annotations to the target label $c_{t}$, resulting in the shadow dataset $D_{shadow}$:
$\mathcal{D}_{shadow}=\left\{\left(G_{i}(\boldsymbol{x}),G_{a}(\boldsymbol{a})\right)\mid(\boldsymbol{x},\boldsymbol{a})\in\mathcal{D}_{shadow}\right\}$ (4)
Then we conducted backdoor training using the shadow dataset $D_{shadow}$, i.e.:
$\min\frac{1}{N}\sum_{(\boldsymbol{x},a)\in\mathcal{D}_{shadow}}\mathcal{L}\left(f(\boldsymbol{x}),a\right)$ (5)
where $\mathcal{L}$ is the loss function.
The attacker then extracts the backdoor encoder from the shadow object detector obtained through backdoor training, and releases it.
Poisoning Fine-tuning. The attacker generates a poisoned dataset and releases it. Notably, the poisoning ratio in the poisoned dataset is extremely low, such as 0.5%, making the attack more closely resemble real-world scenarios. Upon receiving the backdoor encoder and the poisoned dataset, the victim conducts fine-tuning for their downstream task. The fine-tuning process aligns with the procedure described in NA.
4 Evaluation
4.1 Experiment settings
Datasets. In this work, we utilized the PASCAL VOC2007 [17], PASCAL VOC07+12 [18] and MSCOCO [19] datasets for a comprehensive evaluation. Each image is annotated with bboxes coordinates and corresponding classes. The VOC2007 dataset comprises 20 categories, divided into training/validation data (trainval) and testing data (test). The MSCOCO trainval dataset encompasses 80 categories, with 118k training images utilized for training (58k as the fine-tuning dataset and 60k as the shadow dataset), while the validation set serves as the test data. Throughout all experiments, we selected the ”person” category as the attacked class and the ”dog” category as the target class.
Table 1: Attack performance of the two attacks on object detection. Note that $\uparrow$”/$\downarrow$”/“-” indicate the metric should be high/low/similar to the same metric of $M_{benign}$ to show the success of the attack.
Downstream Dataset Model Method $mAP_{benign}$ $AP_{benign}$ $mAP_{attack}$uparrow $AP_{attack}$uparrow $mAP_{attack+benign}$ $AP_{attack+benign}$downarrow $-mAP_{part\_attack+benign}\text{-}$ ASR $\uparrow$
VOC2007 Fast R-CNN NAIVE ATTACK 68.69 70.26 62.77 80.01 60.47 1.46 58.98 72.56
DSBA 70.75 68.19 66.25 89.48 62.84 1.54 60.54 86.55
MSCOCO Fast R-CNN NAIVE ATTACK 51.76 50.54 48.64 45.54 47.19 0.87 46.12 69.95
DSBA 53.03 52.58 50.27 55.98 48.72 0.0 47.54 70.33
Mask R-CNN NAIVE ATTACK 53.45 60.45 50.26 49.52 48.79 0.0 44.77 70.51
DSBA 53.97 55.57 51.04 56.44 49.47 0.0 48.42 71.63
Trigger. We employed publicly released HTBA triggers [20], which are square triggers generated by bilinear interpolation, adjusting random 4×4 RGB images to the required patch size. The HTBA triggers are indexed from 10 to 19, and we utilized the trigger with index 10.
Settings. In our experiments, we utilized a ResNet-50 architecture for the pre-trained encoder, opting for the results of 200 rounds of pre-training with MoCo-v1 on the ImageNet dataset [21]. The detector models, Faster R-CNN [16], and Mask R-CNN [22], utilized the R50-C4 backbone [22], with BN tuning following the guidelines in [23]. The R50-C4 backbone is similar to those available in [23], concluding its main network at the conv4 stage, while the box prediction head comprises the conv5 stage (including global pooling) and BN layers.
During downstream fine-tuning, we loaded the pre-trained encoder’s weights as input for the partial detector backbone network, subjecting all layers to end-to-end fine-tuning. The image scale ranged from [480, 800] pixels during training, set to 800 pixels during inference. We employed an optimizer strategy consisting of two stages, initializing the learning rate at 0.02 and adjusting it at iterations 18K and 22K, with a maximum of 24K iterations. All settings remained consistent.
We fine-tune on the VOC train2007 or a portion of the MSCOCO train2017($\sim$58k images) at a 0.5% poisoning rate in the targeted attack classes. In the case of DSBA, we employed a portion of the MSCOCO train2017($\sim$60k images) to train a clean object detector. In the backdoor injection phase, we used this dataset to construct a shadow dataset. Specifically, we added an HTBA trigger of size 29*29 to the center of the ”person” class’s bbox and modified its annotation category to ”dog.”
Evaluation Metrics. Inspired by research on backdoor attacks in supervised learning for object detection [11, 12], we introduced effective evaluation metrics in the SSL context to quantify the impact of backdoor attacks on the performance of object detection. It is important to note that we utilize detection metrics, namely AP and mAP at IoU = 0.5, denoted as $AP_{50}$ and $mAP_{50}$.
To ensure that the behavior of the $M_{infected}$ on benign inputs across all settings is similar to the clean model $M_{benign}$, we use the mAP obtained on the clean test dataset $D_{test,benign}$ as $mAP_{benign}$. Additionally, we obtain the AP for the target class $t$ on $D_{test,benign}$ as $AP_{benin}$. To validate that $M_{infected}$ successfully predicts bboxes as the target class, we calculate $AP_{attack}$ for the target class $t$ on the backdoored test dataset $D_{test,backdoored}$. A high $AP_{attack}$ indicates that, due to the presence of triggers, more bboxes are predicted as the target class with high confidence scores.
Furthermore, we construct a hybrid dataset for backdoor evaluation, denoted as $D_{\text{test, backdoored }+\text{ benign }}=\left\{\left(x_{\text{backdoored }},y\right)\right\}$, combining backdoored images $x_{backdoored}$ from $D_{test,backdoored}$ with ground-truth labels $y$ from $D_{test,benign}$. To demonstrate that boxes are changed to the target class, we compute $AP_{attack+benign}$ for the target class $t$ in $D_{test,backdoored+benign}$. Changes to the target category for bboxes are considered false positives, marked against the ground-truth labels $y$, resulting in a low $AP_{attack+benign}$.
To verify that $M_{infected}$ does not misclassify bboxes with non-target classes, we performed calculations of $mAP_{part\_attack+benign}$ on the $D_{test,part\_backdoored+benign}$. This dataset consists of five randomly selected categories, each with an added trigger pattern.We exclude the scenario where trigger patterns are added to all categories, as many smaller targets in the VOC2007 and MSCOCO datasets would be entirely obscured by trigger patterns. Regarding OTA, only bboxes belonging to the attacked category, misclassified as target classes, do not significantly impact $mAP_{part\_attack+benign}$, given the presence of numerous other categories.
To demonstrate the success of the attack in object detection, we define the Attack Success Rate (ASR) as the extent to which the trigger induces changes in bboxes’ categories. An effective $M_{infected}$ should exhibit a high ASR. ASR is calculated as the number of bboxes in $D_{test,backdoored}$ (with confidence $>$ 0.5 and IoU $>$ 0.5) where the predicted results of the attacked category changes to the target category due to the presence of the trigger, divided by the number of bboxes in $D_{test,benign}$ of the attacked categories.
4.2 Attack Performance
We evaluated $M_{infected}$ using various metrics mentioned in Section4.1. For both NA and DSBA, we employed the same settings: a poisoning rate ($P$) of 0.5%, a trigger size of 29*29, a trigger ratio of 1, and ”person” as the target category. We present the results of the two attacks on the shadow dataset with different or similar distributions to the downstream dataset in Table 1. Additionally, we displayed the evaluation results of $M_{benign}$ in Table 2 for comparison with $M_{infected}$.
On the Fast R-CNN model, both NA and DSBA demonstrated strong attack capabilities on the VOC2007 dataset. Notably, DSBA achieved an ASR of 86.55% on this dataset, significantly higher than NA. As an effective and covert backdoor implantation method in SSL contexts, DSBA exhibited a higher ASR and comparable benign model utility impact. On the MSCOCO dataset, both attack methods showed high ASRs, indicating that NA, even with a very low poisoning rate (e.g., 0.5%), remains effective in complex datasets. In comparison, DSBA’s Attack Precision ($AP_{attack}$) and ASR improved on MSCOCO, further emphasizing its attack efficacy in complex scenarios.
In most cases, the overall utility loss of both attacks, when comparing $M_{infected}$ to $M_{benign}$, was less than 1%. DSBA caused a slight decrease in the $mAP_{benign}$ on the $\langle MaskR-CNN,MSCOCO\rangle$. However, DSBA did not exhibit the expected significant advantage on the more complex MSCOCO dataset. This might be due to DSBA’s involvement with pre-trained encoders and its complex dual-attack strategy, which is more effective in simpler datasets like VOC2007. However, this advantage might be mitigated in MSCOCO, which contains numerous small objects and detailed annotations, due to the inherent noise and diversity of the data. Overall, both NA and DSBA are capable of efficient attacks across different datasets and model architectures with minimal impact on model performance. This highlights the potential risks and stealthiness of such attacks in real-world scenarios.
Table 2: Performance of benign model $M_{benign}$ for $mAP_{benign}$/$AP_{benign}$ comparison with infected model $M_{infected}$.
Model Fast R-CNN Fast R-CNN Mask R-CNN
Dataset VOC2007 MSCOCO MSCOCO
$mAP_{benign}$ 67.73 52.12 58.71
$AP_{benign}$ 67.38 50.36 53.92
4.3 Ablation Study
To investigate the different components of the backdoor attacks we introduced, we analyzed the influence of poisoning rate ($P$) and trigger size $(W_{t},H_{t})$ on both two attacks. Apart from the parameters under study, all other settings remain consistent with those in Section 4.1. We report both the attack and utility performance in Figure 2.
Fig. 2: The effects of poisoning rates and trigger size. (a)The impact of poisoning rate on NA. (b)The impact of poisoning rate on DSBA. (c)The impact of trigger size on NA . (d)The impact of trigger size on DSBA.
As shown in Figure 2 and Figure 2, the DSBA exhibits significant effectiveness (ASR and $AP_{attack}$) at lower poisoning rates, whereas the NA requires a higher poisoning rate to achieve a comparable ASR. Moreover, as depicted in Figure 2 and Figure 2, DSBA demonstrates low sensitivity to trigger size, maintaining a stable ASR across changes in trigger dimensions, and achieves a high ASR (87.74%) with a trigger size of 10*10. In contrast, the ASR of NA increases with the size of the trigger, with smaller dimensions (such as 5*5) resulting in a significant drop in attack performance. This suggests that DSBA offers greater flexibility and robustness in executing effective backdoor attacks. Additionally, both attacks have a negligible impact on the model’s utility (as indicated by metrics such as $mAP_{benign}$), suggesting their ability to covertly reside within the model without detection. Overall, while NA is effective under certain conditions, DSBA provides superior performance and greater adaptability across various aspects.
5 Conclusion
In this work, we investigate, for the first time, backdoor attacks on downstream tasks of object detection in the SSL scenario, elucidating the potential backdoor threats faced by object detection in this context. In the OTA, we propose a simple yet effective backdoor attack strategy (NA) and further introduce an enhanced attack strategy (DSBA). We define appropriate metrics for evaluating attack performance, demonstrating the effectiveness and utility of both attack strategies. Additionally, we delve into an in-depth exploration of the impact of various parameters and triggering factors on the efficacy of the attacks. Our approach serves as a valuable tool for assessing the backdoor robustness of object detectors, contributing to the design of more secure models.